OpenStudy (anonymous):

Multiply or divide as indicated. 2x/4x+2 * 14x+7/6 B. 7/6 C. x/6 D.7x/6

OpenStudy (anonymous):

@Vlery @neer2890 @Kendra1998

OpenStudy (anonymous):

@ShailKumar

OpenStudy (vlery):

@hartnn can help you out with this one because it 's confusing me

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

@neer2890

OpenStudy (anonymous):

@Vlery can you tag some ppl to help me please

OpenStudy (vlery):

well @hartnn is really good that's the only person in math I know

OpenStudy (anonymous):

well they havent came to the question yet :(

OpenStudy (neer2890):

\[\frac{ 2x }{ 4x+2 }*\frac{ 14x+7 }{ 6 }=\frac{ 2x }{ 2(2x+1) }*\frac{ 7(2x+1) }{ 6 }=\frac{ 7x }{ 6 }\]